19. You purchase a 5-year and 10%-coupon rate bond with face amount of $1,000 at par value. One year later, market interest rate increases from 10% to 20%, what is the rate of capital gain for the bond?

(A) -40.3%.

(B) -29.5%.

(C) -25.9%.

(D) -30.3%.

(E) -15.9%

20.Refer to Q19, what is the rate of return for the bond?

(A) -40.3%.

(B) -38.4%.

(C) -25.9%.

(D) -30.3%.

(E) -15.9%.
Answer #1
19] Price of the bond one year later = 1000/1.2^4+100*(1.2^4-1)/(0.2*1.2^4) = $           741.13
Rate of capital gain = 741.13/1000-1 = -25.9%
Answer: Option [C]
20] Rate of return for the bond = (100+741.13-1000)/1000 = -15.9%
Answer: Option [E]